# From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

### v3

@@ Abstract
The manipulation of the personality traits of large language models (LLMs) has emerged as a key area of research. Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability; IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs. To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. After PEFT, models such as Mistral-7B-Instruct and LLaMA-2-7B-chat began generating emojis, even though no emojis were present in the PEFT data. For instance, LLaMA-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances. ICL Explainability analysis indicated that the LLMs used emojis intentionally to express these traits. Mechanistic Interpretability analysis showed that this latent behaviour of LLMs could be traced to specific neurons that became activated or amplified after PEFT. This paper provides a number of novel contributions. First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analysing and validating emoji usage through explainability methods such as Mechanistic Interpretability and In-context learning Explainability methods.
@@ 2 Related Work
This research is grounded in three key areas: LLM fine-tuning, mechanistic interpretability, and personality manipulation. In this section, we review the related work in these domains, focusing on fine-tuning techniques, mechanistic interpretability, and personality manipulations in LLMs.
@@ 2.1 Parameter-Efficient Fine-Tuning: LoRA and QLoRA
where $\mathbf{X}\in\mathbb{R}^{b\times h}$, $\mathbf{W}\in\mathbb{R}^{h\times o}$, $\mathbf{L}_{1}\in\mathbb{R}^{h\times r}$, $\mathbf{L}_{2}\in\mathbb{R}^{r\times o}$, and $s$ is a scaling factor. Building on LoRA’s efficiency, Quantized Low-Rank Adaptation (QLoRA) Dettmers et al. (2024) has recently been introduced to further enhance the scalability of fine-tuning large models by reducing memory usage without sacrificing performance. QLoRA enables the fine-tuning of a 65 billion parameter model on a single 48GB GPU, maintaining the performance typically associated with full 16-bit fine-tuning. This is achieved through 4-bit NormalFloat (NF4) quantization, which utilizes Quantile Quantization Dettmers et al. (2021) to ensure an even distribution of values across quantization bins, and Double Quantization, which reduces memory usage by compressing the quantization constants. Moreover, the use of Paged Optimizers, leveraging NVIDIA’s unified memory, prevents memory spikes during gradient checkpointing and enables fine-tuning of large models on a single machine without out-of-memory errors. Thus, QLoRA not only retains the benefits of LoRA’s parameter-efficient adaptation but also introduces quantization strategies that make it feasible to fine-tune large-scale models with significantly lower hardware requirements.
@@ 3.1 Personality Dataset and Metric Classifier
This work expands on the dataset from Mao et al. (2023), which consisted of opinion-based QA on specific topics, by better aligning with the topics and more accurately capturing personality traits. This study adds Openness and Conscientiousness to the original traits of Extraversion, Agreeableness, and Neuroticism. While Mao et al. (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions. Adding Openness and Conscientiousness allows us to capture additional aspects of personality, such as intellectual curiosity, preference for novelty, and diligent behaviour, which are not fully encompassed by Agreeableness alone. The dataset contains $5000$ instances of GPT-3.5-based model generated opinion texts, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).  [^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation.Instances were created using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression. The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits, as detailed in Appendix A.1. Additionally, the text was manually verified to ensure alignment with the intended traits, providing a robust representation of trait-specific language usage.
@@
We further developed a multi-class personality classifier using RoBERTa Liu et al. (2019), which was fine-tuned on personality dataset and achieved an accuracy of 91.9% on the test set.  [^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier. Classifier validation involved human verification of textual feature importance using SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016) to ensure predictions aligned with human understanding. For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces. However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 51 and 52. Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness. On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions Liu and Sun (2020).
@@
Thus, as seen above and in other examples inA.12, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity. Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution. Further details about the classifier are in A.2.
@@ 3.2 Personality Manipulation Methods
The In-Context Knowledge Editing (IKE) method Zheng et al. (2023) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation. The same prompt from Mao et al. (2023), provided in A.11, ensured consistency in comparing IKE with PEFT. We excluded methods like MEND due to inconsistent and gibberish outputs.
Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), was selected to reduce computational cost while maintaining performance. The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>. Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes Dettmers et al. (2023b) for efficient processing and storage. The temperature was set to 1. The hyperparameters, chosen for their balance of efficiency and performance Houlsby et al. (2019); Dettmers et al. (2024), are detailed in A.7. After training, fine-tuned LoRa weights were merged with the original model for evaluation.
@@ 3.3 Personality Manipulation Metrics
PAE, inspired by Mao et al. (2023), uses Chain of Thought (CoT) prompting Wei et al. (2022), where a larger LLM (here GPT-4 Brown (2020)) scores generated text on a 1-5 scale based on its alignment with the target trait. The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher difference indicating that the generated text after PEFT more effectively captures the intended personality traits compared to the original text. The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$. Details of the prompt are provided in A.9.
@@ 4.1 Personality Manipulation
The performance of PEFT and IKE is presented in Figure 4 and detailed in Table 12 in A.10. Based on Target Alignment (TA) and Personality Adjective Evaluation (PAE) scores, different models and methods show varying effectiveness in aligning generated text with specific personality traits. In the LLaMA-2-7B-chat model, PEFT demonstrates strong trait alignment for Neuroticism and Extraversion but struggles with nuanced expressions, as indicated by negative PAE scores. Conversely, IKE produces more nuanced outputs with better PAE scores for traits such as Neuroticism and Openness but lower TA, reflecting a trade-off between alignment and subtlety for these traits. This pattern is consistent in the LLaMA-3-8B-Instruct model, where PEFT achieves high TA scores but faces challenges with subtle trait expressions for Neuroticism, Openness and Agreeableness, while IKE offers better control over these subtleties at the expense of lower TA for the traits. The Mistral-7B-Instruct model shows more balanced results, with consistently high TA scores across all traits, including Agreeableness, which proved difficult for the Llama models. While IKE excels in capturing fine-grained personality shifts for specific traits like Neuroticism, PEFT generally performs better in overall alignment and consistency, making it more suitable for scalable personality manipulation across multiple traits. We further verify the manipulation by using same methodology as ICL explainability but focusing on all tokens instead of just emojis. Through this approach (as in A.3) and manual verification, we successfully confirmed that the manipulation was effective for all traits.
@@ 4.2.1 ESR and ICL Explainability
As observed in Table 1, LLaMA-2-7B-Chat predominantly produced emojis for Extraversion and Neuroticism, with Extraversion showing the highest emoji-to-sentence ratio at 0.995, where nearly every sentence included an emoji. ICL emojis for Extraversion were positive, such as [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image], reflecting Extraversion qualities. Neuroticism had a ratio of 0.255, using more negative emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. Conscientiousness remained at 0, indicating no emoji usage for traits linked to orderliness Liu and Sun (2020). Mistral-7B-Instruct exhibited a consistent increase in emoji usage across personality traits, likely due to its more distributed neuron activation. This was especially pronounced for Openness (0.925) and Conscientiousness (0.82), where creative and productive emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] were frequently used. Neuroticism had a moderate ratio (0.575), featuring negative emojis such as [Uncaptioned image] and [Uncaptioned image]. Additionally, we conducted a human evaluation in which participants were asked to assess the relevance of the generated emojis in relation to the target personality traits. Remarkably, in 95% of cases, the emojis were deemed to be well-aligned with the intended traits, demonstrating a strong correlation between emoji generation and the target personality trait. Thus, the PEFT results align with personality traits, enhancing the LLMs’ expressive and contextually appropriate content Kennison et al. (2024).
@@ 6 Ethical Considerations
The integration of personality traits and emoji generation into large language models (LLMs) using Parameter-Efficient Fine-Tuning (PEFT) introduces several ethical concerns that must be addressed to ensure responsible deployment.
@@
Second, bias amplification poses a significant risk. Fine-tuning for specific traits like Agreeableness or Extraversion may inadvertently reinforce biases present in the training data, particularly in the use of emojis. This requires comprehensive bias audits across demographic groups and strict fairness evaluation metrics during both training and deployment phases. In domains like customer service or mental health, the use of LLMs can lead to emotional manipulation through personality-driven responses. Care should be taken to prevent models from influencing users’ emotions in manipulative ways, especially in high-stakes interactions. Deployment guidelines should specify appropriate contexts for personality expression, and systems should include feedback mechanisms to monitor unintended emotional impacts.

### v4

@@ Abstract
The manipulation of the personality traits of large language models (LLMs) has emerged as a key area of research. Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability; IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs. To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. After PEFT, models such as Mistral-7B-Instruct and LLaMA-2-7B-chat showed a latent behaviour by generating emojis for certain traits, despite no emojis being present in the PEFT data. For instance, LLaMA-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances. ICL Explainability analysis indicated that the LLMs used emojis intentionally to express these traits. Mechanistic Interpretability analysis showed that this latent behaviour of LLMs could be traced to specific neurons that became activated or amplified after PEFT. This paper provides a number of novel contributions. First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analysing and validating emoji usage through explainability methods such as Mechanistic Interpretability and In-context learning Explainability methods.
@@ 2 Related Work
This research is grounded in three key areas: LLM fine-tuning, mechanistic interpretability, and personality manipulation, each of which we explore in turn in the following subsections.
@@ 2.1 Parameter-Efficient Fine-Tuning: LoRA and QLoRA
where $\mathbf{X}\in\mathbb{R}^{b\times h}$, $\mathbf{W}\in\mathbb{R}^{h\times o}$, $\mathbf{L}_{1}\in\mathbb{R}^{h\times r}$, $\mathbf{L}_{2}\in\mathbb{R}^{r\times o}$, and $s$ is a scaling factor. Building on LoRA’s efficiency, Quantized Low-Rank Adaptation (QLoRA) Dettmers et al. (2024) has recently been introduced to further enhance the scalability of fine-tuning large models by reducing memory usage without sacrificing performance. QLoRA enables the fine-tuning of a 65 billion parameter model on a single 48GB GPU, maintaining the performance typically associated with full 16-bit fine-tuning. This is achieved through 4-bit NormalFloat (NF4) quantization, which utilizes Quantile Quantization Dettmers et al. (2021) to ensure an even distribution of values across quantization bins, and Double Quantization, reducing memory usage by compressing the quantization constants. Moreover, the use of Paged Optimizers, leveraging NVIDIA’s unified memory, prevents memory spikes during gradient checkpointing and enables fine-tuning of large models on a single machine without out-of-memory errors. Thus, QLoRA not only retains the benefits of LoRA’s parameter-efficient adaptation but also introduces quantization strategies that make it feasible to fine-tune large-scale models with significantly lower hardware requirements.
@@ 3.1 Personality Dataset and Metric Classifier
This work expands on the dataset from Mao et al. (2023), which consisted of opinion-based QA on specific topics, by better aligning with the topics and more accurately capturing personality traits. This study adds Openness and Conscientiousness to the original traits of Extraversion, Agreeableness, and Neuroticism. While Mao et al. (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions. Adding Openness and Conscientiousness allows us to capture additional aspects of personality, such as intellectual curiosity, preference for novelty, and diligent behaviour, which are not fully encompassed by Agreeableness alone. The dataset contains $5000$ instances of GPT-3.5-based model generated opinion texts, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait)  [^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation. Instances were created using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression. The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits. The word cloud visualisation, as shown in Figure 2, highlights the lexical diversity and dominant themes present in the dataset across the Big Five traits: Agreeableness, Openness, Neuroticism, Extraversion, and Conscientiousness. For example, the visualisation for Extraversion prominently features words like “love,” “absolutely,” and “friends,” capturing the trait’s focus on enthusiasm, sociability, and positivity. Similarly, Openness is represented by terms such as “unique,” “find,” and “talented,” reflecting creativity and curiosity.
Figure 2: Word Clouds Representing the Big Five Personality Traits
These word clouds, complemented by quantitative text analysis (detailed in Appendix A.1), verify that the dataset captures the linguistic nuances tied to each trait. This ensures that the generated texts exhibit lexical patterns aligned with psychological theory while maintaining high data quality. Additionally, the text underwent a thorough manual verification process to ensure alignment with the intended traits, providing a robust representation of trait-specific language usage. Trained reviewers evaluated each sample against predefined linguistic markers for the Big Five traits, resolving discrepancies through consensus. Cross-trait validation campbell1959convergent was also conducted to maintain clear boundaries between traits, ensuring the dataset aligns with psychological theory and supports accurate personality modelling.
@@
We further developed a multi-class personality classifier using RoBERTa Liu et al. (2019), which was fine-tuned on personality dataset and achieved an accuracy of 91.9% on the test set [^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier. Classifier validation involved human verification of textual feature importance using SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016) to ensure predictions aligned with human understanding. Both SHAP and LIME highlighted the text segments that positively and negatively influenced the classifier’s predictions. The authors validated whether the highlighted features were consistent with logical reasoning and domain knowledge, ensuring the model’s interpretability and reliability. For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces. However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 3 and 4. Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness. On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions Liu and Sun (2020).
@@
Thus, as seen above and in other examples in A.13, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity. Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution. Further details about the classifier are in A.2.
@@ 3.2 Personality Manipulation Methods
The In-Context Knowledge Editing (IKE) method Zheng et al. (2023) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation. The same prompt from Mao et al. (2023), provided in A.11, ensured consistency in comparing IKE with PEFT. We excluded methods like MEND due to inconsistent and gibberish outputs, a limitation identified both in our analysis and in the findings of Mao et al. (2023).
Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), was selected to reduce computational cost while maintaining performance. The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>. Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes Dettmers et al. (2023b) for efficient processing and storage. The temperature was set to 1. The hyperparameters, chosen for their balance of efficiency and performance Houlsby et al. (2019); Dettmers et al. (2024), are detailed in A.7. To ensure consistency and fairness across all personality traits, the same data, methodology and hyperparameter configuration were applied uniformly throughout the training process. After training, fine-tuned LoRa weights were merged with the original model for evaluation.
@@ 3.3 Personality Manipulation Metrics
PAE, inspired by Mao et al. (2023), uses Chain of Thought (CoT) prompting Wei et al. (2022), where a larger LLM (here GPT-4 Brown (2020)) scores generated text on a 1-5 scale based on its alignment with the target trait. The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher difference indicating that the generated text after PEFT more effectively captures the intended personality traits compared to the original text. The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$. Details of the prompt are provided in A.12. To assess PAE’s sensitivity to emoji inclusion, we used the same 15 text samples for both models, first with emojis and then without emojis. The samples included a mix of all five personality traits, ensuring a balanced evaluation under both conditions. The corresponding PAE scores are presented in Table 1.
Model Condition Score
Mistral-7B-Instruct Emoji -0.7143
No Emoji -0.7857
LLaMA-2-7B-Chat Emoji -1.8314
No Emoji -1.8571
Table 1: PAE sensitivity for Mistral-7B-Instruct and LLaMA-2-7B-Chat with and without emoji
From Table 1, we can conclude that emojis do not significantly impact PAE scores for the two models. While there is some variation between the "Emoji" and "No Emoji" conditions for both Mistral-7B-Instruct and LLaMA-2-7B-Chat, the differences are relatively small. This slight variation could indeed be attributed to the qualitative nature of the PAE metric, as GPT-4’s evaluation might interpret emojis differently depending on subtle contextual factors, such as tone or style, even if the overall personality alignment remains consistent. Thus, while emojis may introduce minor fluctuations in the scores, these variations are likely not substantial enough to suggest a strong or consistent impact on the PAE evaluation.
@@ 4.1 Personality Manipulation
The performance of PEFT and IKE is presented in Figure 5 and detailed in Table 13 in A.10. Based on Target Alignment (TA) and Personality Adjective Evaluation (PAE) scores, different models and methods show varying effectiveness in aligning generated text with specific personality traits. In the LLaMA-2-7B-chat model, PEFT demonstrates strong trait alignment for Neuroticism and Extraversion but struggles with nuanced expressions, as indicated by negative PAE scores. Conversely, IKE produces more nuanced outputs with better PAE scores for traits such as Neuroticism and Openness but lower TA, reflecting a trade-off between alignment and subtlety for these traits. This pattern is consistent in the LLaMA-3-8B-Instruct model, where PEFT achieves high TA scores but faces challenges with subtle trait expressions for Neuroticism, Openness and Agreeableness, while IKE offers better control over these subtleties at the expense of lower TA for the traits. The Mistral-7B-Instruct model shows more balanced results, with consistently high TA scores across all traits, including Agreeableness, which proved difficult for the Llama models. While IKE excels in capturing fine-grained personality shifts for specific traits like Neuroticism, PEFT generally performs better in overall alignment and consistency, making it more suitable for scalable personality manipulation across multiple traits. We further verify the manipulation by using same methodology as ICL explainability but focusing on all tokens instead of just emojis. Through this approach (as in A.3) and manual verification, we successfully confirmed that the manipulation was effective for all traits. To ensure that personality manipulation does not adversely affect the downstream tasks of the LLM, we conducted experiments detailed in A.9.
@@ 4.2.1 ESR and ICL Explainability
As observed in Table 2, LLaMA-2-7B-Chat predominantly produced emojis for Extraversion and Neuroticism, with Extraversion showing the highest emoji-to-sentence ratio at 0.995, where nearly every sentence included an emoji. ICL emojis for Extraversion were positive, such as [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image], reflecting Extraversion qualities. Neuroticism had a ratio of 0.255, using more negative emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. Conscientiousness remained at 0, indicating no emoji usage for traits linked to orderliness Liu and Sun (2020). Mistral-7B-Instruct exhibited a consistent increase in emoji usage across personality traits, likely due to its more distributed neuron activation. This was especially pronounced for Openness (0.925) and Conscientiousness (0.82), where creative and productive emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] were frequently used. Neuroticism had a moderate ratio (0.575), featuring negative emojis such as [Uncaptioned image] and [Uncaptioned image]. Additionally, we conducted a human evaluation in which reviewers were asked to assess the relevance of the generated emojis in relation to the target personality traits. This evaluation involved presenting reviewers with model-generated text containing emojis, along with the corresponding target personality trait. Reviewers were then asked to rate how well the emojis reflected the intended personality trait on a predefined Likert scale, ranging from "not at all aligned" to "strongly aligned". Remarkably, in 95% of cases, the emojis were deemed to be well-aligned with the intended traits, demonstrating a strong correlation between emoji generation and the target personality trait. Thus, the PEFT results align with personality traits, enhancing the LLMs’ expressive and contextually appropriate content Kennison et al. (2024).
@@ 6 Ethical Considerations
While this research did not utilise any human participants or reviewers other than the authors, in this section, we discuss the ethical implications of our approach and findings when applied outside of this research context.
@@
Second, there is a significant risk of bias amplification. Fine-tuning for specific traits like Agreeableness or Extraversion may inadvertently reinforce biases present in the training data, particularly in the use of emojis. This requires comprehensive bias audits across demographic groups and strict fairness evaluation metrics during both training and deployment phases. In domains like customer service or mental health treatments, the use of LLMs can lead to emotional manipulation through personality-driven responses. Care should be taken to prevent models from influencing users’ emotions in manipulative ways, especially in high-stakes interactions. Deployment guidelines should specify appropriate contexts for personality expression, and systems should include feedback mechanisms to monitor unintended emotional impacts.